OpenStudy (techarena):

Need explanation on how to find the product matrix of two 3*3 matrix.

OpenStudy (anonymous):

When two matrices are 3x3 Then op matrix is also 3x3. Multiply each row element of 1st matrix with each column element of second matrix and add the values. Example : Row is 2 3 4 and column is 1 2 3 Multiply and add = 2*1 + 2 * 4 + 2*3 = 16 = 3*1+3*2+3*3 = 19 = 4*1 + 4*2 + 4*3 = 24 Hence op matrix first row is 16 19 24. You can use this https://www.easycalculation.com/matrix/learn-matrix-multiplication.php https://www.easycalculation.com/matrix/matrix-multiplication.php

OpenStudy (thomas5267):

Look at this. I learned matrix multiplication using this website. http://www.purplemath.com/modules/mtrxmult.htm